# Boosting Mixed-Initiative Co-Creativity in Game Design: A Tutorial

### v1

@@ 4.2.4. RL Brush
Initial setting: In this case, the conceptual space is not restricted to any specific game genre, and the tool can prove useful for creating different kinds of 2D assets beyond maps. The computational agent offers a wide range of templates that the human agent can choose from, but the human agent also has the option to add their own custom templates. Despite the computer offering several alternatives for the input that will serve as the basis to generate the image, it is ultimately the human designer’s decision on which one to use. The co-creative process can be initiated by the human agent by manually creating a desired area on the canvas and then requesting the computational agent to continue filling in the remaining space. Alternatively, the computational agent can start generating content from a blank canvas, but only upon request from the user. As such, although both agents can contribute to the initial concept, it is the human agent who retains more control and determines the specific problem they want to address.
Initiative: The human agent holds more initiative throughout the creative process. The extent of contributions from the computational agent can vary significantly depending on how much the human designer allows it to intervene. Although the computational agent is practically capable of generating the entire creative artifact, it can only do so with permission from the human agent. Hence, the computational agent is not proactive. Furthermore, the system aims to generate an image similar to the initial input and aligned with the tile properties adjusted by the user. It does not actively seek to create content that is novel and surprising, thus being less creative than the human co-creator.
Evaluation: When the human agent is manually editing the canvas, the computational agent assesses the tile or set of tiles that the human partner intends to insert and may reject the edit if it deems it unfitting. However, this evaluation performed by the system focuses on the accuracy with respect to the example input provided for the generation, rather than the creativity of the human agent’s contribution. In contrast, the human agent continuously evaluates the quality of the content created by both themself and the computational agent, revisiting previous stages and making changes if deemed necessary. Thus, the evaluation process is primarily conducted by the human agent.
Final decision: The human designer is the one who decides when they are satisfied with the created map or texture and ready to export it.
Task assignment: The agents are involved in the same task of creating a 2D visual asset, such as a game level or texture. Both the human agent and the computational agent can directly edit the entire area of the bitmap, and their contributions strongly influence each other. However, there are some actions that the human agent can perform but the computational agent cannot. Specifically, the human agent can delete content generated by the system and manipulate tile properties.
Intervention pace: The co-creators have well-defined shifts. When the human agent is editing the image, the computational agent waits and only intervenes upon the human partner’s request. Despite this, the computational agent can generate content immediately, enabling rapid interaction between the agents. On the other hand, for the human designer to manually edit the canvas, they must switch screens, inevitably slowing down the intervention pace. The amount of content produced by each agent during their turn varies according to how the human agent chooses to interact with the computational agent. The computational agent can generate anything from a single tile to a complete bitmap, depending on the extent to which the human partner allows.
Explainability: There is little explainability between the human agent and the computational agent. All the user knows is that the system attempts to create content similar to the template used as an example and according to the tile properties defined by the user. Additionally, the system’s display of an animation of the generative process that the user can analyze step by step allows for a better understanding of the creative process of the computational agent. Nevertheless, the human agent has no other means of gaining further insight into the computational agent’s decisions, nor does the computational agent have a way to comprehend the creative objectives and preferences of the human agent.
4.2.6. Pitako
@@
Initial setting: At the onset of the creative process, the system consistently presents a default background sprite and three default suggestions. However, the human agent can immediately add a new sprite, triggering further suggestions from the computational agent. Consequently, while the human designer may be influenced and even choose to accept the initial suggestions, they retain control over the addition of the first sprite that truly begins to shape the game.It is worth noting that the sprites available for the human agent to add are limited to the behaviors and images found within the system’s catalog. While the catalog is extensive, it still imposes creative limitations on the human partner. Considering all these factors, we posit that the initial setting criterion leans more toward the computer end of the spectrum.
Initiative: There are several actions that the human agent can perform. They can manually add sprites, accept sprite suggestions, manipulate various properties of each sprite, define interactions, accept interaction suggestions, position sprites on the game map, and define rules for game completion. On the other hand, the computational agent can only provide recommendations for sprites and interactions. Regarding sprite suggestions, the system autonomously provides them whenever the human designer adds a new sprite to the set. However, there comes a point when the sprite set for the developing game becomes extensive, and the system can no longer generate new suggestions based on that sprite set. As for interaction suggestions, these are presented only upon explicit user request. Thus, the human agent possesses considerably greater initiative than the computer agent. The sprite suggestion is the only part of the process where the computer is proactive, but even in this case, its contribution is limited in creativity as it relies on associations from existing games.
Evaluation: The evaluation is primarily conducted by the human designer, who assesses their contributions and the game they are designing, as well as the system’s recommendations. The computational agent solely self-evaluates its suggestions in terms of “confidence” and communicates this assessment to the human partner. A high confidence value indicates that the suggested element is common in existing games, while a low confidence value indicates that it is less common, making the suggestion more novel. However, the computational agent does not conduct that same evaluation for the sprites and interactions added by the user.
Final decision: The human agent is the one who determines when they are satisfied with the game, the chosen sprites, and the implemented dynamics.
Task assignment: The agents share some tasks in the creative process, but not all. Specifically, both are involved in creating sprites, mechanics, and dynamics. The human agent can directly add these elements to the game, whereas the computational agent only presents them as recommendations. Furthermore, only certain actions by the human creator directly impact the next contribution from the computer (e.g., adding a sprite with a behavior already present in the set does not trigger new suggestions). In the case of sprite placement on the level grid and the definition of termination rules, the computational agent does not contribute to these tasks.
Intervention pace: The pace and spontaneity of co-creators’ interventions depend on whether they are working on creating sprites or interactions. In the case of sprites, the turns are short, with the computational agent typically providing suggestions after each intervention by the human partner. The content leap is not too large since both parties focus on one sprite at a time. However, creating a sprite involves not only defining its graphical representation but also its behavior, which, in turn, may require specifying several properties. Thus, we can neither consider the content leap to be small in this context. In the case of interactions, the system does not respond spontaneously to the contributions of the human creator, providing recommendations only upon request and solely based on the sprite set. Consequently, the user receives no feedback as they add interactions to the game.
Explainability: Explainability is almost non-existent. The computational agent provides a small hint regarding the novelty of its recommendations by displaying the confidence level, leaving it to the discretion of the human partner whether they want to create a game that closely resembles existing ones or opt for a set of sprites and interactions that is more distinct. However, it does not provide any further information about its creative process, and likewise, the human agent cannot justify their contributions either.
4.2.7. Story Designer
@@
Initial setting: In any case, the co-creative process begins with the system consistently presenting the same base graph, as well as immediately providing a set of suggestions. Consequently, the human agent is inevitably prompted to shape the narrative structure around this initial content. While it is possible for the human designer to erase the entire graph and start anew, the initial setting presented by the computational agent is most likely to influence on their creative process. Regarding the conceptual space, although the nodes encapsulate a high degree of abstraction that allows for numerous creative possibilities to elaborate the narrative afterward, they are nonetheless rooted in tropes. Hence, they remain considerably constrained by the narrative conventions commonly prevalent in existing games as identified by the system.
Initiative: During the co-creative process, the human agent retains the capability to directly edit the narrative graph, whereas the computational agent provides suggestions prompted by human interventions without explicit request. Although the computational agent contributes autonomously, with a focus on presenting novel and interesting content, it remains unable to manipulate the main graph. Furthermore, each suggestion only becomes visible if the user opts to inspect it. As a result, the system might offer multiple suggestions that the user may never come across. For these reasons, the human agent has the stronger initiative.
Evaluation: Both parties evaluate the creative artifacts. The human agent self-assesses the content they create and the suggestions presented by the computational agent. Conversely, the computational agent evaluates both the main graph and the suggestions it generates in terms of feasibility, while also considering other metrics such as interestingness and coherence. Additionally, the computational agent discloses its assessment of the main narrative graph, the hovered suggestion, and the edited suggestion by displaying fitness, interestingness, and coherence scores for each. However, regardless of the computational agent’s evaluation of the quality and creativity of the artifacts, only the human agent holds the authority to accept suggestions if they prefer them over the current graph.
Final decision: It is the human agent who determines when the design of the narrative structure is complete.
Task assignment: Both agents are engaged in the design of narrative structures and can produce the same type of content - narrative graphs. However, the human agent can manipulate the main graph by editing nodes and connections individually, whereas the computational agent contributes by presenting suggestions for complete graphs. The interventions made by the human agent directly influence the subsequent input of the computational agent, but the reverse may or may not occur.
Intervention pace: The system is not swift enough to generate suggestions in real-time as the user makes alterations, often resulting in delayed responses. Moreover, to view the suggestions generated by the computational agent, the human partner needs to inspect them one by one, thereby slowing down the intervention pace. As for the extent of the content leap, it varies from suggestion to suggestion. For instance, suggestions with a higher value in the “step” dimension will entail a more significant content leap.
Explainability: Similar to the Sentient Sketchbook, the system displays the fitness, interestingness, and coherence scores of suggestions in relation to the main narrative graph. This provides some insight into the specific optimizations certain suggestions offer and why they might be preferable. However, this level of explainability remains insufficient, and there is no means for the human agent to reciprocally provide justification.
4.2.8. 1001 Nights
@@
Initial setting: The human agent initiates the creative process by being the first to write, setting the starting point from which the story will unfold. Nevertheless, the computational agent still exerts some influence on the initial setting. In addition to continuing the player’s narrative, the AI model is prompted to focus on an ancient Middle Eastern story related to the keywords representing all valid weapons. This, coupled with the game’s own narrative setting and the goal of generating these weapons, steers the story toward this particular literary genre.
Initiative: Both the human and the computational agent exhibit a strong initiative as they contribute equally by writing chunks of the narrative alternately. They are both attentive and adapt to each other’s input in a creative manner. The ability to directly craft the story through writing engages both participants deeply in the creative process. The only difference lies in the human co-creator having a limited number of characters for their writing turn, which allows the system to generate more extensive content.
Evaluation: Both agents assess what the other has written, choosing to adapt to their colleague’s storyline and further develop their idea or change the direction or topic of the narrative. However, the human agent can assess the story as a whole. In contrast, the computational agent can only recall the last five inputs from both itself and the human co-creator. While they respond to each other’s input thoughtfully, neither agent can critique the creativity and quality of the contributions from their partner.
Final decision: Since the narrative creation serves as a means to generate game mechanics, there is no aim or incentive for the story to reach a conclusion (unlike when storytelling itself is the primary goal). Nonetheless, it is the human agent who determines when they are ready to engage in battle with the King, thus concluding the co-creative storytelling session.
Task assignment: Both the human and the computational agent are performing the same task and in the same manner. The creative work of one agent directly affects what their partner will write next.
Intervention pace: In this approach, the agents have precise turns where each one contributes a portion of the story alternately. Although the turns are relatively short, we consider the content leap to be moderate, especially on the side of the computational agent, as it can generate more content in its turn. While the human agent may write one or two sentences, the computational agent can write an entire paragraph uninterrupted. The intervention pace would be faster if the agents could react while their counterpart is writing.
Explainability: Since the game’s objective is for the player to prompt the King to mention keywords that can be converted into weapons, the player may sometimes become puzzled when the system’s response lacks any of these keywords. In such cases, the system provides the player with a hint that they should tell a story more centered around swords and magic. However, this explanation remains identical for every instance in which the King fails to mention any of the items and proves ineffective if the player is already narrating a story focused on these topics. Furthermore, this explanation only vaguely alludes to the underlying process of the computational agent concerning the use of target words, providing no further insight into its overall creative process in story writing. Consequently, explainability is almost absent.
4.2.9. Germinate
@@
Initial setting: The Germinate tool is tailored for a style of gameplay with simple mechanics, yet it still allows for the expression of specific rhetorical intent. Specifically, the games created with this tool can only be played with the mouse, lack a game map, and offer limited variety in the types of dynamics possible between entities and resources. Consequently, part of the conceptual space is constrained by the computational agent. However, the meaning assigned to the game, its goals, rules, and interactions are yet to be defined. Typically, it is the human agent who establishes the initial rule set. The system provides some example cards to get things started, but it is up to the user to determine which entities, resources, relationships, and triggers they want to add. The human agent can also request the computational agent to generate a random card, but ultimately, the initial concept of the game is mostly determined by the human agent.
@@
Initiative: Once the human agent specifies their intention by defining the cards, the computational agent generates a batch of complete and playable games, each with its established rules and dynamics. If the human co-creator has allowed it beforehand, the computational agent can also generate entities and resources for each game beyond those defined by the human partner. Throughout the co-creative process, the system recurrently generates more batches of games for the user to explore. After the human agent makes changes to the properties and requests the computational agent to generate a new set of games, the current game list is cleared and the computational agent starts generating new games based on the updated settings. In this scenario, we consider the computational agent to have more initiative as it can continuously create various complete games, whereas the human agent can only define cards and cannot directly edit any game. Initially, the computational agent may not be proactive, but once prompted by the human agent, it autonomously generates games. Furthermore, despite the games generated by the system being constrained by the user’s expressed intent, the system can typically add a fair amount of content to the game beyond what the user established. Therefore, the computational agent still retains a significant degree of creative freedom.
@@
Evaluation: The evaluation is primarily conducted by the human agent. While the computational agent can generate numerous games, it is the human agent who explores them to identify any with desirable characteristics. Based on this assessment, the human agent may modify their intent, leading to the discarding of all previous games to generate a new list. The only assessment made by the system is to check for contradictions in the cards defined by the user that would prevent game generation, without specifying the nature of these contradictions. Importantly, the computational agent does not provide any input or judgment on the meaning and creativity of the rhetorical intent of the human co-creator.
Final decision: The human agent is the one who decides whether they are satisfied with any of the games and when the creative process comes to an end.

### v2

@@ 4.2.1. Initial Setting
miWFC (mW): In this case, the conceptual space is not restricted to any specific game genre, and the tool can prove useful for creating different kinds of 2D assets beyond maps. The computational agent offers a wide range of templates that the human agent can choose from, but the human agent also has the option to add their own custom templates. Despite the computer offering several alternatives for the input that will serve as the basis to generate the image, it is ultimately the human designer’s decision on which one to use. The co-creative process can be initiated by the human agent by manually creating a desired area on the canvas and then requesting the computational agent to continue filling in the remaining space. Alternatively, the computational agent can start generating content from a blank canvas, but only upon request from the user. As such, although both agents can contribute to the initial concept, it is the human agent who retains more control and determines the specific problem they want to address.
@@
Pitako (P): At the onset of the creative process, the system always presents a default background sprite and three default suggestions. However, the human agent can immediately add a new sprite, triggering further suggestions from the system. Consequently, while the human designer may be influenced and even choose to accept the initial suggestions, they retain control over the addition of the first sprite that truly begins to shape the game. It is worth noting that the sprites available for the human agent to add are limited to the behaviors and images found within the system’s catalog. While the catalog is extensive, it still imposes creative limitations on the human partner. Considering all these factors, we posit that the initial setting criterion leans more toward the computer end of the spectrum.
@@
Story Designer (SD): The co-creative process begins with the system consistently presenting the same base graph and immediately providing a set of suggestions. Consequently, the human agent is inevitably prompted to shape the narrative structure around this initial content. While it is possible for the human designer to erase the entire graph and start anew, the initial setting presented by the computational agent is most likely to influence their creative process. Regarding the conceptual space, although the nodes encapsulate a high degree of abstraction that allows for numerous creative possibilities to elaborate the narrative afterward, they are nonetheless rooted in tropes. Hence, they remain considerably constrained by the narrative conventions commonly prevalent in existing games.
@@
1001 Nights (1N): The human agent initiates the creative process by being the first to write, setting the starting point from which the story will unfold. Nevertheless, the computational agent still exerts some influence on the initial setting. In addition to continuing the player’s narrative, the AI model is prompted to focus on an ancient Middle Eastern story related to the keywords representing all valid weapons. This, coupled with the game’s own narrative setting and the goal of generating these weapons, steers the story toward this particular literary genre.
@@
Germinate (G): The tool is tailored for a style of gameplay with simple mechanics, yet it still allows for the expression of specific rhetorical intent. Specifically, the games created with Germinate can only be played with the mouse, lack a game map, and offer limited variety in the types of dynamics possible between entities and resources. Consequently, part of the conceptual space is constrained by the computational agent. However, the meaning assigned to the game, its goals, rules, and interactions are yet to be defined. Typically, it is the human agent who establishes the initial rule set. The system provides some example cards to get things started, but it is up to the user to determine which entities, resources, relationships, and triggers they want to add. The human agent can also request the computational agent to generate a random card, but ultimately, the initial concept of the game is mostly determined by the human agent.
Figure 12. Overview of the ranking of the Initial Setting for each work.
4.2.2. Initiative
AI Spaceship Generator (SG): The computational agent can only intervene by generating several spaceship suggestions, while the user can only select one option and request new generation rounds. The user is the one who decides whether there will be a new round of spaceship generation and from what point (whether from the selected spaceship, a random spaceship, or resetting the population). However, the user has no way of directly modifying the artifacts or clearly indicating their preferred characteristics, restricting their creativity. The computational agent only generates more spaceships at the behest of its partner but can create their entire structure, exerting significantly greater control over the visual appearance of the artifacts. Thus, we consider the system to have more initiative.
Sprite Editor for Pixel Art Characters (SE): The human agent has the stronger initiative. They can directly edit any pose view, including those in the suggestion slots. By being able to draw and promptly modify any artifact, the human agent is granted more creative freedom. Meanwhile, the computer can only intervene upon user request, lacking proactivity to make independent contributions. It provides suggestions but cannot manipulate the artifacts directly. Moreover, the computational agent always tries to match the design provided by the human partner, avoiding the suggestion of images that are conceptually different from the source. Thus, its input lacks creativity.
Sentient Sketchbook (SS): The human agent contributes by directly editing the map sketch and selecting suggestions. The computational agent intervenes by consistently suggesting alternative map designs but cannot directly edit the main sketch. Both agents are highly proactive and engage multiple times throughout the creative process. The human partner can continuously add tiles and analyze suggestions, while the computer willingly responds with new creative input after each contribution made by its counterpart. Thus, both agents are deeply involved in the process and attentive to each other’s feedback. Still, the user slightly stands out due to the ability to directly manipulate the map sketch.
RL Brush (RB): Similar to the Sentient Sketchbook, the human agent directly impacts the creative artifact throughout the process by being able to manually edit the level grid or choose a suggestion to replace it. The computational agent provides suggestions whenever the human partner makes a modification, but cannot directly edit the current level. This approach has the additional characteristic that the user can control the performance of the system, granting them the power to limit the quality and scope of its suggestions. Consequently, both agents proactively contribute creative input, but the human designer retains more control over the process.
miWFC (mW): The human agent holds more initiative throughout the creative process. The extent of contributions from the computational agent can vary significantly depending on how much the human designer allows it to intervene. Although the computational agent is practically capable of generating the entire creative artifact, it can only do so with permission from the human agent. Hence, the computational agent is not proactive. Furthermore, the system aims to generate an image similar to the initial input and aligned with the tile properties adjusted by the user. It does not actively seek to create content that is novel and surprising, thus being less creative than the human co-creator.
Pitako (P): There are several actions that the human agent can perform. They can manually add sprites, accept sprite suggestions, manipulate various sprite properties, define interactions, accept interaction suggestions, position sprites on the game map, and define rules for game completion. On the other hand, the computational agent can only provide recommendations for sprites and interactions. Regarding sprite suggestions, the system autonomously provides them whenever the human designer adds a new sprite to the set. However, there comes a point when the sprite set becomes extensive, and the system can no longer generate new suggestions based on that set. As for interaction suggestions, these are presented only upon explicit user request. Thus, the human agent has considerably greater initiative. On top of that, the computer’s contribution is limited in creativity as it relies on associations from existing games.
Story Designer (SD): During the co-creative process, the human agent retains the capability to directly edit the narrative graph, whereas the computational agent provides suggestions prompted by human interventions without explicit request. Although the computational agent contributes autonomously, with a focus on presenting novel and interesting content, it remains unable to manipulate the main graph. Furthermore, each suggestion only becomes visible if the user opts to inspect it. As a result, the system might offer multiple suggestions that the user may never come across. For these reasons, the human agent has the stronger initiative.
1001 Nights (1N): Both the human and the computational agent exhibit a strong initiative as they contribute equally by writing chunks of the narrative alternately. They are both attentive and adapt to each other’s input in a creative manner. The only difference lies in the human co-creator having a limited number of characters for their writing turn, which allows the system to generate more extensive content.
@@ 4.2.2. Initiative
Germinate (G): Once the human agent specifies their intention by defining the cards, the system generates a batch of playable games, each with its established rules and dynamics. If the human co-creator has allowed it beforehand, the system can also generate entities and resources for each game beyond those defined by the human partner. Throughout the co-creative process, the system recurrently generates more batches of games for the user to explore. After the human agent makes changes to the properties and requests the system to generate a new set of games, the current game list is cleared, and the system starts generating new games based on the updated settings. We consider the computational agent to have more initiative as it can continuously create various complete games, whereas the human agent can only define cards and cannot directly edit any game. Furthermore, despite the games generated by the system being constrained by the user’s expressed intent, the system can typically add a fair amount of content to the game beyond what the user established. Therefore, the computational agent still retains a significant degree of creative freedom.
Figure 13. Overview of the ranking of the Initiative for each work.
4.2.3. Evaluation
AI Spaceship Generator (SG): The human agent evaluates the artifacts by selecting the one they prefer to use as a basis for a new iteration. However, the system can only predict the reasons that lead the human agent to prefer a particular spaceship over others. That is, the evaluation criteria of the human agent are assumed. In theory, the system favors user preferences and adjusts the fitness function accordingly, but this is not guaranteed in practice.
Sprite Editor for Pixel Art Characters (SE): The system provides suggestions that align with the input from the user without having an opinion about what is being drawn. The evaluation process is entirely carried out by the human agent, who determines whether to accept or ignore suggestions and modify the content if they deem it necessary.
Sentient Sketchbook (SS): Both evaluate the content being created. The human agent self-assesses the content they are creating, and the map suggestions of the system, possibly choosing to replace the current sketch or to draw inspiration from certain features of the suggestions. On the other hand, the computer tests the playability of the main map sketch and evaluates it based on gameplay properties related to game pace and player balance. It does the same to generate map designs and choose the best ones to present to the user. Additionally, half of the map designs generated by the system are evaluated in terms of novelty, and the system selects the most distinct maps to include in the suggestions. There is, however, a significant difference in the evaluation process carried out by the two agents: the system can only disclose its opinion to the human partner regarding gameplay characteristics. It cannot express whether it finds a particular map creative or not. Furthermore, regardless of what the computational agent informs the user about its evaluation, only the user can accept or reject suggestions. Therefore, the human agent has more evaluative power.
RL Brush (RB): The evaluation is mostly done by the human agent. They not only self-assess the content they create but also evaluate the suggestions provided by the computational agent, deciding whether any of them are good enough to be accepted. As for the computational agent, the internal evaluation it performs to generate suggestions remains unclear. We lack insight into the considerations that the AI models of the system have regarding the modifications they propose and how these suggestions may enhance the current level. Even if the system does conduct some form of evaluation, it cannot express it. Therefore, it is the human agent who guides the level design process, retaining complete control over the desired outcome for the creative artifact.
miWFC (mW): When the human agent is manually editing the canvas, the computational agent assesses the tile or set of tiles that the human partner intends to insert and may reject the edit if it deems it unfitting. However, this evaluation performed by the system focuses on the accuracy with respect to the example input provided for the generation, rather than the creativity of the human agent’s contribution. In contrast, the human agent continuously evaluates the quality of the content created by both themself and the computational agent, revisiting previous stages and making changes if deemed necessary. Thus, the evaluation process is primarily conducted by the human agent.
Pitako (P): The evaluation is primarily conducted by the human designer, who assesses their contributions and the game they are designing, as well as the system’s recommendations. The computational agent solely self-evaluates its suggestions in terms of “confidence” and communicates this assessment to the human partner. A high confidence value indicates that the suggested element is common in existing games, while a low confidence value indicates that it is less common, making the suggestion more novel. However, the computational agent does not conduct that same evaluation for the sprites and interactions added by the user.
Story Designer (SD): Both parties evaluate the creative artifacts. The human agent self-assesses the content they create and the suggestions presented by the computational agent. Conversely, the computational agent evaluates both the main graph and the suggestions it generates in terms of feasibility, while also considering other metrics such as interestingness and coherence. Additionally, the computational agent discloses its assessment of the main narrative graph, the hovered suggestion, and the selected suggestion by displaying fitness, interestingness, and coherence scores for each. However, regardless of the computational agent’s evaluation of the quality and creativity of the artifacts, only the human agent holds the authority to accept suggestions if they prefer them over the current graph.
1001 Nights (1N): Both agents assess what the other has written, choosing to adapt to their colleague’s storyline and further develop their idea or change the direction or topic of the narrative. However, the human agent can assess the story as a whole. In contrast, the computational agent can only recall the last five inputs from both itself and the human co-creator. While they respond to each other’s input thoughtfully, neither agent can critique the creativity and quality of the contributions from their partner.
@@ 4.2.3. Evaluation
Germinate (G): The evaluation is primarily conducted by the human agent. While the system can generate numerous games, it is the user who explores them to identify any with desirable characteristics. Based on this assessment, the user may modify their intent, leading to the discarding of all previous games to generate a new list. The only assessment made by the system is to check for contradictions in the cards defined by the user that would prevent game generation, without specifying the nature of these contradictions. Importantly, the computational agent does not provide any input or judgment on the meaning and creativity of the rhetorical intent of the human co-creator.
Figure 14. Overview of the ranking of the Evaluation for each work.
4.2.4. Final decision
AI Spaceship Generator (SG), Sprite Editor for Pixel Art Characters (SE), RL Brush (RB), miWFC (mW), Pitako (P), Story Designer (SD), Germinate (G): The human agent is the one who decides when they are satisfied with an artifact (spaceship, character sprites, level map, texture, dynamics, narrative structure, or game, depending on the tool) and ready to end the creative process.
Sentient Sketchbook (SS): The human agent determines when they believe the map sketch is finalized and ready to be exported. Subsequently, the computational agent generates a detailed view of the map. The human partner cannot directly edit the detailed version of the map, but can choose their preferred option from the viewing modes provided by the computer or request the generation of a new view. Thus, the human has the final say over the structure of the level but has less control over its final look.
1001 Nights (1N): Since the narrative creation serves as a means to generate game mechanics, there is no aim for the story to reach a conclusion (unlike when storytelling itself is the primary goal). Nonetheless, it is the human agent who determines when they are ready to engage in battle with the King, thus concluding the co-creative storytelling session.
Figure 15. Overview of the ranking of the Final decision for each work.
4.2.5. Task assignment
AI Spaceship Generator (SG): Even though both agents are working on creating spaceship designs, they do not work in the same way. The computational agent intervenes by generating the entire structure of the spaceships, while the human agent can only contribute by selecting a spaceship they prefer among those presented to them. This selection has no disruptive impact on the computer’s subsequent intervention.
Sprite Editor for Pixel Art Characters (SE): Both agents are engaged in the creation of sprites for a pixel art character. The human agent has the ability to draw the character from scratch, directly manipulate any pose view, and select each image that will serve as input for a new suggestion. On the other hand, the computational agent can generate sprites for the character but can only present them as suggestions when prompted. In short, we can assert that both agents are able to generate content within the same domain, although not to the same degree.
Sentient Sketchbook (SS): The agents are collaborating on the shared task of creating map designs for a strategy game. The input from the human agent directly affects each subsequent intervention of the computational agent, whereas the contributions of the computational agent may or may not impact the next user intervention. Although the computer cannot edit the main sketch and select suggestions, it can generate map suggestions autonomously. Thus, both agents willingly construct map designs.
RL Brush (RB): Both agents are working on the shared task of designing a level for the Sokoban game. The contributions made by the human agent directly influence the suggestions presented by the computational agent. Conversely, the creative input from the computational agent may or may not affect the creative process carried out by the human co-creator. Lastly, even though the system cannot directly edit the grid of the current level, both agents contribute by adding or altering tiles.
miWFC (mW): The agents are involved in the same task of creating a 2D visual asset, such as a game level or texture. Both the human agent and the computational agent can directly edit the entire area of the bitmap, and their contributions strongly influence each other. However, there are some actions that the human agent can perform but the system cannot. Specifically, the human agent can delete content generated by the system and manipulate tile properties.
Pitako (P): The agents share some tasks in the creative process, but not all. Specifically, both are involved in creating sprites, mechanics, and dynamics. The human agent can directly add these elements to the game, whereas the computational agent only presents them as recommendations. Furthermore, only certain actions by the human creator directly impact the next contribution from the computer (e.g., adding a sprite with a behavior already present in the set does not trigger new suggestions). In the case of sprite placement on the level grid and the definition of termination rules, the computational agent does not contribute to these tasks.
Story Designer (SD): Both agents are engaged in the design of narrative structures and can produce the same type of content - narrative graphs. However, the human agent can manipulate the main graph by editing nodes and connections individually, whereas the computational agent contributes by presenting suggestions for complete graphs. The interventions made by the human agent directly influence the subsequent input of the computational agent, but the reverse may or may not occur.
1001 Nights (1N): Both the human and the computational agent are performing the same task and in the same manner. The creative work of one agent directly affects what their partner will write next.